Rivoli Inc. hired you as a consultant to help estimate its cost of capital. You have been provided with the following data: D0 = $0.80; P0 = $77.50; and g = 8.00% (constant). Based on the DCF approach, what is the cost of equity from retained earnings?

Group of answer choices

8.66%

7.20%

10.12%

9.11%

10.30%

Malholtra Inc. is considering a project that has the following cash flow and WACC data. What is the project's MIRR? Note that a project's MIRR can be less than the WACC (and even negative), in which case it will be rejected.

WACC: 10.00%
Year           0           1       2          3         4    
Cash flows -$850 $300 $320 $340 $360

Group of answer choices

14.08%

15.65%

17.21%

18.94%

20.83%

Last month, Lloyd's Systems analyzed the project whose cash flows are shown below. However, before the decision to accept or reject the project, the Federal Reserve took actions that changed interest rates and therefore the firm's WACC. The Fed's action did not affect the forecasted cash flows. By how much did the change in the WACC affect the project's forecasted NPV? Note that a project's projected NPV can be negative, in which case it should be rejected.

Old WACC: 10.00% New WACC 12.50%
Year 0 1 2 3
Cash flows -$1,000 $410 $410 $410

Group of answer choices

-$43.26

-$39.80

-$48.02

-$47.15

-$44.12
Answer #1

1)

Cost of equity =D1/P0+g

=(0.80*1.08)/77.50+8%

=9.11%

2)

Year 1 2 3 4 CF $(850.00) $ 300.00 $ 320.00 $ 340.00 $ 360.00 15.65% MIRR

Year ICF 0 -850 1 300 320 340 4 360 MIRR] =MIRR(B2:B6,10% 10%)!

3)

Year Old WACC New WACC 0 $(1,000.00) $(1,000.00) 1 $ 410.00 $ 410.00 2 $ 410.00 $ 410.00 3 $ 410.00 $ 410.00 Diff. $ 19.61 $

Year Old WACC New WACC 01-1000 T-1000 1 1410 |410 2 410 410 3 410 410 Diff. Ī=NPV(10%,B3:B5)+B2 I=NPV(12.5%,C3:05)+C2 =C6-B6